OpenStudy (anonymous):

shane drew a square with each side equal to n units: part A: shane increased the length and width of of the square by m units. what will be the change in the area of the original square? part B: shane decreased the length and width of the of the square by m units. what will be the change in the area of the square? part C: shane increased the length of the square by m units and decreased its width by m units. what will be the change in the area of the square?

OpenStudy (ranga):

Original square has side of length n units. Original Area = n^2 A) New square has side of length (n+m) units. New Area = (n+m)^2 Change in area is: (n+m)^2 - n^2 = n^2 + 2nm + m^2 - n^2 = m^2 + 2nm or m(m+2nm) Do B) and C) in a similar way.

OpenStudy (anonymous):

wait how do you do part B? do you subtract?

OpenStudy (anonymous):

@ranga

OpenStudy (anonymous):

@Zarkon help?

OpenStudy (anonymous):

does anybody know how to solve part b?????????????????????????????????

OpenStudy (skullpatrol):

yes

OpenStudy (skullpatrol):